OpenStudy (anonymous):

is (2,4) a solution to my equation y=x-2 ?

OpenStudy (anonymous):

Ok try plugging it into the equation and see if you get the same thing So, input either x or y into the equation and see what you get. (x,y) so x = 2, y=4 if you plug in x, you should get 4 as an answer. try it.

OpenStudy (anonymous):

2-2=4 ?

OpenStudy (anonymous):

y = 2 -2 y= 0 4 does not equal 0. so it's not a solution

OpenStudy (anonymous):

yes, you set it up right.

OpenStudy (anonymous):

and (-3,-9) y = 3x isn't a solution also?

OpenStudy (anonymous):

-9 = 3x -9/3 -3 = x Yes it is a solution

OpenStudy (anonymous):

or y = 3(-3) y = -9